# DAWN: Domain-Adaptive Weakly Supervised Nuclei Segmentation via Cross-Task Interactions

### v1

@@ Abstract
Weakly supervised segmentation methods have gained significant attention due to their ability to reduce the reliance on costly pixel-level annotations during model training. However, the current weakly supervised nuclei segmentation approaches typically follow a two-stage pseudo-label generation and network training process. The performance of the nuclei segmentation heavily relies on the quality of the generated pseudo-labels, thereby limiting its effectiveness. This paper introduces a novel domain-adaptive weakly supervised nuclei segmentation framework using cross-task interaction strategies to overcome the challenge of pseudo-label generation. Specifically, we utilize weakly annotated data to train an auxiliary detection task, which assists the domain adaptation of the segmentation network. To enhance the efficiency of domain adaptation, we design a consistent feature constraint module integrating prior knowledge from the source domain. Furthermore, we develop pseudo-label optimization and interactive training methods to improve the domain transfer capability. To validate the effectiveness of our proposed method, we conduct extensive comparative and ablation experiments on six datasets. The results demonstrate the superiority of our approach over existing weakly supervised approaches. Remarkably, our method achieves comparable or even better performance than fully supervised methods. Our code will be released in https://github.com/zhangye-zoe/DAWN.
@@ I Introduction
Nuclei segmentation [1, 2]plays a vital role in computational pathology, serving as a fundamental step for various nuclei-related tasks. Accurate segmentation enables advanced analysis, such as nuclei classification [3, 4, 5], nuclei counting [6, 7] and nuclei tracing [8, 9, 10]. However, developing supervised nuclei segmentation models in whole slide images (WSIs) is challenging due to the extensive requirement of precise pixel annotation in these images, which consist of gigapixels. At the same time, manual labeling of nucleus boundaries is time-consuming and demands domain expertise. To alleviate the manual burden, a weakly supervised segmentation algorithm is a promising solution, which leverages the inherent image structure and context to achieve segmentation, facilitating more efficient and accurate nuclei analysis.
@@
Weakly supervised point annotation methods have recently been widely adopted for nuclei segmentation [11, 12, 13], which can decrease manual consumption by a large margin. However, due to inadequate supervision information, these methods generally perform a pre-processing step to generate masks for model training, as shown in Fig. 1 (a). Subsequently, these generated pseudo-labels are utilized to guide the training of the segmentation network. However, this two-stage training approach still faces two significant challenges. Firstly, the pseudo-labels generated through pre-processing often exhibit a coarse granularity and lack precise representation of nuclear morphology. For example, the Voronoi method [11] utilizes distance measurements to differentiate overlapping nuclei, which results in the disruption of boundary morphology, as depicted in Fig. 2 (b). The level set method (LSM) [14] employs color features to simulate nuclear contours. However, it fails to distinguish tissues with similar colors, as illustrated in Fig. 2 (c). Secondly, these two-stage methods still have problems handling the suboptimal pseudo-labels. Although some single-stage optimization methods have been proposed to enhance segmentation performance, such as utilizing boundary mining techniques [15] to improve the quality of pseudo-labels and employing conditional random fields (CRF) [16] for post-processing to rectify segmentation results, these localized improvement methods fail to adapt to the entire training process, resulting in segmentation outcomes that remain unsatisfactory. Therefore, adopting a new training strategy is necessary to mitigate the impact of pseudo-label generation on nuclei segmentation results.
@@
Recent research has demonstrated the potential of domain adaptation in nuclei instance segmentation [17, 18, 19]. They leverage prior information from the source domain and excavate domain-invariant features from the source data to achieve cross-domain task transfer, resulting in remarkable performance. Therefore, it is reasonable to believe that domain-adaptive methods can enhance instance segmentation performance in weakly supervised scenarios. On the one hand, domain adaptation methods do not explicitly generate pseudo-labels for supervising network training [20, 21]. Instead, they enhance the network’s performance in the target domain through domain transfer strategies. It can mitigate the impact of inaccurate pseudo-labels during model training. On the other hand, compared to unsupervised domain adaptation methods, weakly supervised domain adaptation has evident superiority for providing additional supervision information. In the context of weak supervision, incorporating an auxiliary task, such as object detection, can leverage the weak annotations to effectively enhance the segmentation network’s ability to achieve domain transfer [22, 23, 24].
@@
Motivated by these observations, this paper proposes a Domain-Adaptive Weakly supervised Nuclei segmentation method (DAWN). By utilizing point annotation data, DAWN incorporates a cross-task interaction strategy at both the feature and output levels, aiming to enhance the model’s domain transfer capability for accurate nuclei segmentation in weak supervision scenarios. Firstly, we propose a consistent feature constraint (CFC) module, which leverages the task correlation between the detection and segmentation networks to preserve prior knowledge of the target data. Secondly, we introduce a task-level supervision information optimization strategy to address the challenge of the domain mismatch problem. By leveraging the strong localization ability of the detection network, we compensate for the output shortcomings of the segmentation network. Lastly, we devise an interactive training method to achieve progressive transfer of the segmentation network. In detail, we establish mutual supervision between the segmentation and detection tasks. This approach enhances the performance of segmentation predictions by leveraging the complementary aspects of both tasks. Extensive experiment results demonstrate the superiority of the proposed framework. Overall, our contributions can be summarized as follows:
@@
  • We propose a novel domain-adaptive weakly supervised nuclei segmentation framework that departs from traditional two-stage weakly supervised methods. Our approach leverages a detection task with point annotations and domain transfer strategies to address the issue of inaccurate pseudo-label generation.
@@
  • We devise a consistent feature constraint module, which utilizes prior information of the source domain to improve the nuclei segmentation performance of target domain.
@@
  • We develop a combined pseudo-label optimization method incorporating the output information from the segmentation and the detection networks. Besides, through an iterative training strategy, this module enhances the domain transfer capability at the task level.
@@
  • We propose an interactive supervised training method that leverages optimized pseudo-labels to simultaneously supervise the training of segmentation and detection networks. This approach facilitates the learning of complementary features between the two networks.
@@
  • We validate our method on six datasets. Extensive experiments indicate that our method outperforms existing weakly supervised methods and achieves similar or even better performance with fully supervised methods.
@@ II Related Work
II-A Fully Supervised Nuclei Segmentation
Nuclei segmentation is an essential workflow, and analysis based on nuclei can provide a basis for downstream tasks of computational pathology. In the past, watersheds [25] or Otsu threshold[26] usually led to over-segmentation or under-segmentation problems. With the successes of U-Net[8] in the cell tracing task, the variations of U-Net are emerging [27, 28, 29] and widely used in the nuclei segmentation. At present, nuclei segmentation methods mainly focus on solving the nuclei overlapping and unclear staining problems. To solve nuclei overlapping, HoverNet[30] calculates the horizontal and vertical gradients to supervise training. Cellpose[31] uses simulated diffusion and spatial gradients to trace instances to which each pixel belongs. CDNet[32] employs the direction difference map to distinguish different instances of overlapping regions. Besides, DCAN [33] adds the contour category to the semantic prediction task to separate instances. To solve the unclear staining problem, some staining separation methods [34, 35] are proposed to extract nuclei staining components. In addition, Sams-net [36] and Triple-Unet [37] use H-channel of staining separation images to enhance the nuclei feature representation ability. Although these methods achieve good progress in nuclei segmentation, they need a large amount of pixel-level manual annotation, which is expensive and labor-intensive. On the contrary, weakly supervised methods provide an effective solution to reduce the reliance of model development on annotations. They allow training models with coarse annotations, reducing time and manual costs.
II-B Weakly Supervised Segmentation
Existing weakly supervised segmentation labels mainly include image-level, bounding-box, and point annotations. Among that, the models using image-level annotations typically rely on the Class Activation Map (CAM) [38] to generate pseudo labels. PRM[39] utilizes peak backpropagation from CAMs to generate highly informative object regions. SC-CAM[40] accelerates the integrity of CAMs using the unsupervised clustering sub-category information. RRM[41] uses the CAM and reliable score evaluation to generate finer pseudo labels. The methods utilizing bounding-box annotations mainly focus on the saliency models. BoxCaseg[42] employs the salient segmentation head to enhance the object localization ability. BoxTeacher[43] uses noise-aware pixel loss and noise-reduced affinity loss to optimize the pseudo labels. Due to the small nuclei morphology, point labels are widely adopted as the supervision information in pathological images. WeakSeg [16] develops Voronoi splitting and clustering to generate pseudo labels. After that, Voronoi is widely used to generate segmentation masks[11, 44, 45, 13]. Besides, [46] establishes a propagating detection map to segment fluorescence images. [12] adopts a co-training strategy to segment phase images. Even though these weakly supervised nuclei segmentation methods have achieved good results, they often lead to poor shape prediction due to the need for more morphological consideration. In this paper, we improve the nuclear morphology through a pre-training model and domain adaptation strategies.
II-C Domain Adaptation in Medical Image Analysis
Pre-training methods [47] can bring convenience to small-sample task training. However, directly using pre-trained models often leads to poor performance due to the distribution differences of samples. Therefore, some domain adaptation methods are proposed to address the problem of domain transfer. [48] uses the image translation method to align the pixel value distribution, reducing the gap between the source domain and the target domain. DIGA[49] develops a distribution adaptation and a semantic adaptation module to enable model adaptation jointly. Inspired by natural scenario methods[50], domain adaptation is also used in medical image analysis[17]. DARCNN[19] implements domain adaptation from natural scene images to hematoxylin-eosin (H&E) images, exploiting domain-invariant and domain-specific features. CyC-PDAM[51] transfers the model from the fluorescence images domain to the H&E images domain. CAPL-Net [52] implements domain adaptation in H&E images from different sources. SAMDA [53] combines the Segment Anything Model (SAM) with nnUNet [54] in the embedding space to achieve a few-shot domain adaptation. However, existing models mainly utilize unsupervised methods for domain adaptation, often leading to poor performance due to the lack of understanding of target domain features. In this paper, we leverage auxiliary tasks to establish connections between the source domain model and the detection task, which promotes the transfer efficiency of the source model to the target domain.
@@ III-A Framework Overview
Detection and segmentation are closely related tasks that share similar semantic information, such as the location and texture of the object, which can be utilized to assist in instance detection or segmentation. In multiple-task learning scenarios, detection and segmentation networks are usually combined to improve the performance of related tasks [55, 56]. At that point, under the domain adaptation background, we design a detection task based on point annotations to assist the segmentation network domain transfer. By enhancing the adaptability of the segmentation network to target domain data, the nuclei segmentation accuracy can be improved in weakly supervised scenarios.
This paper establishes cross-task interaction strategies at both the feature and task levels. The feature interaction is designed to preserve the prior segmentation information related to the target domain. At the same time, the task interaction is used to compensate for the shortcomings of the segmentation output through combined pseudo-label optimization and interactive supervision methods. Overall, the flowchart of our proposed DAWN is depicted in Fig. 3.
We firstly pre-train a segmentation network in source domain as shown in Fig. 3 (a). To overcome the domain shift problem caused by data distribution variations, we design a consistent feature constraint (CFC) module and a combined pseudo-label optimization (CPL) module as shown in Fig. 3 (c)-(d). In addition, to promote domain transfer ability and efficiency, we employ an interactive supervision training (IST) method. According to these constraints, our training loss function is formulated as follows:
$Loss=L_{det}+\alpha L_{fea}+\beta L_{dyn},$ (1)
where $L_{det}$ is the loss of detection network, $L_{fea}$ is the consistent feature constraint loss, and $L_{dyn}$ is the dynamic supervision loss. $\alpha$ and $\beta$ are used to balance the weight of each loss. It should be noted that $L_{det}$ corresponds to a fully-supervised learning task.
@@ III-B Consistent Feature Constraint
On source domain data $D_{S}$, we firstly pre-train a segmentation network as shown in Fig. 3 (a) and Stage 1 of Algorithm 1. We adopt HoverNet [30] as a segmentation network in this step, and its supervision information consists of three components: foreground probability $p_{s}$, horizontal distance $h_{s,x}$, and vertical distance $h_{s,y}$. Using the pixel-level annotation $y_{s}^{f}$, the supervision images can be obtained through pre-processing, as shown in Fig. 3 (a). In this step, the loss function is formulated as follows:
$\begin{split}L_{pre}&=CE(\hat{p}_{s},p_{s})+Dice(\hat{p}_{s},p_{s})\\ &+MSE(\hat{h}_{s,x},h_{s,x})+MSE(\hat{h}_{s,y},h_{s,y}),\end{split}$ (2)
where $\hat{p}_{s}$, $\hat{h}_{s,x}$, and $\hat{h}_{s,y}$ represent the foreground probability prediction, horizontal distance prediction, and vertical distance prediction. CE, Dice, and MSE represent cross-entropy loss, dice loss, and mean square error loss.
In the following domain adaptation process, we jointly train the segmentation and detection networks, and the process is shown in Fig. 3 (c) and Stage 2 of Algorithm 1. Firstly, the well-trained segmentation network is utilized to initialize the target domain segmentation model. Then, we map the output feature $f_{seg}$ of encoder into $f_{seg}^{c}$ by a convolution layer and a fully connection layer:
@@
where $FC$ and $Conv$ represent fully connected and convolution layers, respectively. Finally, the output feature $f_{seg}^{c}$ is used to compute the consistent feature constraint. It should be noted that the detection network is trained in the target domain, and the segmentation network is initialized in the source domain. Hence, the CFC can force the segmentation network to gradually close to the target domain.
Considering single central pixel of nuclei just represents nuclei position while ignoring nuclei texture, it is hard to train an effective detection network using point label $y_{p}^{t}$ [57, 58]. Thus, we employ extended Gaussian distribution [16] to model nuclei morphology. The extended Gaussian encoding process is shown in Fig. 3 (b) and can be formulated as follows:
@@
where $D_{i}$ represents the distance between pixel $i$ and nearest point annotation, $r_{i}$ represents initialized nuclei radius, the region between $r_{1}$ and $r_{2}$ is background, $\sigma$ is the discrete degree of Gaussian distribution. In $M_{g,i}$, the pixel corresponding to $-1$ does not participate in the training. Employing the generated encoding, a supervised detection model is trained as follows:
$L_{det}=\frac{1}{|\Omega|}\sum_{i\in\Omega}w_{i}(\hat{q}_{i}-M_{g,i})^{2},$ (5)
where $\Omega$ represents the pixel set participating training, i.e. $\Omega=\{i|0<D_{i}\leq r_{2}\}$; $w_{i}$ is the training weight of pixel $i$. In order to pay more attention to the foreground, we set $w_{i}$ equals $10$ when $D_{i}<r_{1}$; otherwise, $w_{i}$ equals $1$. The extended Gaussian encoding not only encodes the tissue texture but can also learn nuclei shape gradually, assisting segmentation network optimization.
Similar to the above segmentation network, the middle layer output $f_{det}$ of the detection network is mapped into $f_{det}^{c}$ to calculate CFC as shown in Fig. 3 (c) and the feature constraint loss $L_{fea}$ is calculated as follows:
@@
where $dim$ represents the dimension of feature encoding. $f_{seg}^{c}$ and $f_{det}^{c}$ are mapped features of segmentation and detection networks. In this process, the CFC module forces the two networks to excavate the essential task correlation features, enhancing domain adaptation efficiency.
@@
11:  Initialize: $r=4$; $e_{t}=20$; Initialize segmentation network with $W_{s}$;
@@
15:        Calculate loss: $L_{det}+\alpha L_{fea}+\beta L_{dyn}$ in Eq. 1 and update network;
@@
17:     Prediction outputs: $\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$;
18:     Update pseudo-label: $M_{pse}$leftarrow CPL($\hat{p}_{t}^{seg}$, $\hat{q}_{t}^{det}$, $y_{t}^{p}$);

### v3

@@ Abstract
Weakly supervised segmentation methods have garnered considerable attention due to their potential to alleviate the need for labor-intensive pixel-level annotations during model training. Traditional weakly supervised nuclei segmentation approaches typically involve a two-stage process: pseudo-label generation followed by network training. The performance of these methods is highly dependent on the quality of the generated pseudo-labels, which can limit their effectiveness. In this paper, we propose a novel domain-adaptive weakly supervised nuclei segmentation framework that addresses the challenge of pseudo-label generation through cross-task interaction strategies. Specifically, our approach leverages weakly annotated data to train an auxiliary detection task, which facilitates domain adaptation of the segmentation network. To improve the efficiency of domain adaptation, we introduce a consistent feature constraint module that integrates prior knowledge from the source domain. Additionally, we develop methods for pseudo-label optimization and interactive training to enhance domain transfer capabilities. We validate the effectiveness of our proposed method through extensive comparative and ablation experiments conducted on six datasets. The results demonstrate that our approach outperforms existing weakly supervised methods and achieves performance comparable to or exceeding that of fully supervised methods. Our code is available at https://github.com/zhangye-zoe/DAWN.
@@ I Introduction
Nuclei segmentation plays a crucial role in computational pathology, serving as a fundamental step for various pathological diagnosis and patient prognosis tasks [1, 2, 3, 4]. Meanwhile, accurate segmentation enables advanced nuclei analysis, including nuclei classification [5, 6, 7], nuclei counting [8, 9], and nuclei tracing [10, 11, 12]. However, developing fully supervised nuclei segmentation models for whole slide images (WSIs) poses challenges because of the extensive requirement of precise pixel-wise annotations in these gigapixel-sized images. To ease the manual annotation burden, the weakly supervised segmentation algorithm provides a promising solution. It leverages weakly supervised annotation information to excavate the morphology and structure of objects, enabling more efficient and accurate nucleic analysis.
@@
Weakly supervised methods based on point annotation have recently been widely adopted for nuclei segmentation [13, 14, 15], which reduce manual consumption by a large margin. Due to inadequate supervision information, these methods generally need a pre-processing step to generate masks. Then, these generated masks are utilized to guide the training of the segmentation network, as shown in 1 (a). However, the two-stage training flow faces two significant challenges. On the one hand, the generated pseudo-labels often are coarse and lack precise representation of nuclear morphology. For example, the Voronoi method [13] utilizes distance measurements to differentiate overlapping nuclei, which results in the disruption of boundary morphology, as depicted in Fig. 2 (b). The level set method (LSM) [16] employs color features to simulate nuclear contours. But, it fails to distinguish cytoplasm and nucleus with similar colors, as illustrated in Fig. 2 (c). On the other hand, these two-stage methods still suffer from the influences of inaccurate pseudo-labels. Although some localized optimization methods have been proposed to enhance segmentation performance, such as boundary mining [17] and conditional random fields (CRF) [18], these localized optimizing methods fail to adapt to the entire training process, resulting in segmentation outcomes that remain unsatisfactory. Therefore, it is necessary to adopt a new training paradigm to improve the nuclear segmentation performance in weakly supervised scenarios.
@@
Recent research has highlighted the potential of domain adaptation (DA) for improving segmentation tasks. By utilizing task-relevant information from the source domain, DA can improve instance segmentation performance in the target domain. However, existing unsupervised domain adaptation methods [19, 20, 21] often struggle with dense segmentation tasks like nuclei segmentation, primarily due to their limited ability to incorporate target domain information effectively. In weakly supervised scenarios, point-annotated images provide valuable prior knowledge regarding the location and distribution of nuclei. This additional information can be leveraged to enhance domain adaptation models. To address this, we propose a novel weakly supervised nuclei segmentation network incorporating domain adaptation techniques. Our approach presents several advantages over previous models. First, unlike traditional weakly supervised segmentation methods that rely heavily on generating pseudo-labels for training [22, 23], our method employs weak annotation to train auxiliary task to boost domain transfer capabilities. This approach minimizes the issues associated with inaccurate pseudo-labels. Second, our weakly supervised segmentation model provides distinct advantages over unsupervised methods by incorporating additional supervisory information. This integration allows us to leverage weak annotation information more effectively, significantly enhancing domain transfer capability [24, 25, 26].
@@
Motivated by these observations, this paper proposes a Domain-Adaptive Weakly supervised Nuclei segmentation method (DAWN). By utilizing point annotation data, DAWN incorporates a cross-task interaction strategy at both the feature and task levels to enhance the model’s domain transfer capability for accurate nuclei segmentation. Firstly, we propose a consistent feature constraint (CFC) module, which leverages the task correlation between the detection and segmentation networks to preserve prior knowledge related to the target data. Secondly, we introduce a task-level supervision information optimization strategy to address the challenge of the domain mismatch problem. By leveraging the strong localization ability of the detection network, we compensate for the output shortcomings of the segmentation network. Lastly, we devise an interactive training method to achieve progressive transfer of the segmentation network. In detail, we establish mutual supervision between the segmentation and detection tasks. This approach enhances the performance of segmentation predictions by leveraging the complementary aspects of both tasks. Extensive experiment results demonstrate the superiority of the proposed framework. Overall, our contributions can be summarized as follows:
@@
  • We propose a novel domain-adaptive weakly supervised nuclei segmentation framework that departs from traditional two-stage weakly supervised methods. Our approach leverages a detection task with point annotations and domain transfer strategies for accurate segmentation.
@@
  • We devise a consistent feature constraint module to retain the prior information from the source domain, which boosts the domain transfer efficiency at the feature level.
@@
  • We develop a combined pseudo-label optimization method incorporating the outputs from segmentation and detection networks. Besides, through an iterative training strategy, this module enhances the domain transfer capability at the task level.
@@
  • We propose an interactive supervised training method that leverages optimized pseudo-labels to supervise the training of segmentation and detection networks simultaneously. This approach facilitates the learning of complementary features between the two networks.
@@
  • We validate our method on six datasets. Extensive experiments indicate that our method outperforms existing weakly supervised methods and achieves equal or even better performance with fully supervised methods.
@@ II Related Work
II-A Nuclei Instance Segmentation
Pathological image analysis based on nuclear segmentation [27, 28, 29, 30, 31] has been widely discussed. Currently, some representative work mainly focuses on the problem of nuclear boundary overlapping. Such as DCAN [32] adds the contour prediction task to the semantic segmentation model to separate instances. HoverNet[33] uses horizontal and vertical gradients to distinguish overlapping instances. Cellpose[34] uses simulated diffusion and spatial gradients to trace the instances to which each pixel belongs. In addition, other methods [35, 36] are also proposed to solve the boundary overlapping problem.
Although fully supervised nuclei segmentation has achieved good progress, it is expensive and labor intensive because it requires a lot of manual annotation at the pixel level. In contrast, weakly supervised methods provide an effective solution to reduce the reliance of model development on annotations. Mainstream weakly supervised nuclear segmentation models follow a two-stage training paradigm. First, point annotation images are used to generate pseudo-labels, and then these generated pseudo-labels are used to supervise network training [13, 37, 38, 39]. However, these two-stage training methods are affected by initialized pseudo-labels, and there is no explicit optimization in the training process, so the model performance is poor. In this paper, we propose a domain adaptation method for weakly supervised nuclear segmentation. Our method uses end-to-end training to avoid the traditional two-stage training paradigm and significantly improves the accuracy of nuclear segmentation in weakly supervised scenarios.
II-B Weakly Supervised Segmentation Methods
Existing weakly supervised segmentation approaches utilize various types of partially labeled annotation to help automatically pixel-level segmentation. These weak labels include image-level, box-level, and point labels. Among these, methods relying on image-level annotations typically use class activation maps (CAMs) [40] to generate pseudo-labels. For instance, PRM [41] employs CAM peak backpropagation to identify highly informative object regions. SC-CAM [42] enhances CAM completeness by incorporating information from unsupervised clustering. RRM [43] combines CAM with a reliable score evaluation to produce more refined pseudo-labels. In addition, several other models [44, 45] are proposed to solve weakly supervised video segmentation tasks. The box-level annotations are often used in conjunction with saliency models. BoxCaseg [46] leverages a salient segmentation head to improve object localization, while BoxTeacher [47] applies noise-aware pixel loss and noise-reduced affinity loss to optimize pseudo-labels. Point labels [48, 31, 49] are widely used in small object segmentation tasks because of their excellent location ability. ASDT [39] propagates labels between image and pixel levels, thereby improving semantic segmentation performance. Despite the success of existing weakly supervised segmentation methods, these approaches often struggle with accurate shape prediction due to insufficient consideration of the morphological characteristics of instances. In this paper, we introduce the domain adaptive method into the weakly supervised segmentation model and learn morphological features with the prior knowledge of external data, thus improving the segmentation performance.
II-C Domain Adaptation Segmentation
Pre-training methods [50] have made it possible to train models with limited samples. However, directly applying pre-trained models to target domains often fails due to distribution gaps between source and target samples. This challenge has led to the development of domain adaptation techniques. For example, Wang et al. [51] use image translation methods to align pixel value distributions, thereby reducing the domain gap. Similarly, DIGA [52] introduces both distribution adaptation and semantic adaptation modules to facilitate joint model adaptation. ADPL [53] improves domain adaptation by constructing an adaptive dual-path network that bridges source and target domains. CyCADA [54] combines spatial alignment of generated images with potential representation alignment to achieve unsupervised domain adaptation. In a different approach, DDMRL [55] addresses object detection by employing two stages domain diversity and multi-domain invariant representation learning. WDASS [56] proposes a weakly-supervised domain adaptive semantic segmentation method that leverages prototypical contrastive learning to align features across domains. Drawing inspiration from natural scene methods [57, 58, 59], domain adaptation has also been applied to medical image analysis [19]. DARCNN [21] adapts models from natural scene images to hematoxylin-eosin (H&E) stained images, capturing both domain-invariant and domain-specific features. CyC-PDAM [60] transfers models between fluorescence images and H&E images, while CAPL-Net [61] focuses on domain adaptation for H&E images from different sources. SAMDA [62] integrates the Segment Anything Model (SAM) with nnUNet [63] in the embedding space to enable few-shot domain adaptation. Despite the progress made by these domain adaptation methods, their performance often suffers due to insufficient understanding of target domain characteristics. In this paper, we propose leveraging auxiliary tasks to bridge the gap between the source domain model and the detection task, enhancing the efficiency of model transfer to the target domain.
@@ III-A Framework Overview
Detection and segmentation are closely related tasks that share semantic information, such as object location and texture, which can be leveraged to enhance both object detection and instance segmentation. In multi-task learning, detection and segmentation networks are often combined to improve the performance of both tasks [64]. In the context of domain adaptation, we design a detection task using point annotations to support the segmentation network’s adaptation to the target domain. By improving the segmentation network’s adaptability to target domain data, this approach enhances nuclei segmentation accuracy, particularly in weakly supervised scenarios.
This paper introduces cross-task interaction strategies at both the feature and task levels. The feature-level interaction is designed to retain prior segmentation information relevant to the target domain. In contrast, task-level interaction compensates for the limitations of segmentation output through a combination of pseudo-label optimization and interactive supervision techniques. The overall flow of our proposed DAWN is illustrated in Fig. 3.
We first pre-train the segmentation network in the source domain, as illustrated in Fig. 3 (a). To address the domain shift caused by variations in the data distribution, we introduce a Consistent Feature Constraint (CFC) module and a Combined Pseudo-Label (CPL) optimization module, as shown in Fig. 3 (c)-(d). Additionally, to enhance the efficiency and effectiveness of domain transfer, we employ an Interactive Supervision Training (IST) method. Based on these components, our overall training loss function is defined as follows:
$Loss=L_{\text{det}}+\alpha L_{\text{fea}}+\beta L_{\text{dyn}},$ (1)
where $L_{\text{det}}$ represents the detection network loss, $L_{\text{fea}}$ is the consistent feature constraint loss, and $L_{\text{dyn}}$ corresponds to the dynamic supervision loss. The parameters $\alpha$ and $\beta$ are used to balance the weight of different loss terms. It should be noted that $L_{\text{det}}$ is associated with a fully supervised learning task.
@@ III-B Consistent Feature Constraint
On source domain data $D_{S}$, we firstly pre-train a segmentation network as shown in Fig. 3 (a) and Stage 1 of Algorithm 1. We adopt HoverNet [33] as a segmentation network in this step, and its supervision information consists of three components: foreground probability $p_{s}$, horizontal distance $h_{s,x}$, and vertical distance $h_{s,y}$. Using the pixel-level annotation $y_{s}^{f}$, the supervision images can be obtained through pre-processing, as shown in Fig. 3 (a). In this step, the loss function [33] is formulated as follows:
$\begin{split}L_{pre}&=\text{CE}(\hat{p}_{s},p_{s})+\text{Dice}(\hat{p}_{s},p_{s})\\ &+\text{MSE}(\hat{h}_{s,x},h_{s,x})+\text{MSE}(\hat{h}_{s,y},h_{s,y})\\ &+\text{MSE}(\nabla\hat{h}_{s,x},\nabla h_{s,x})+\text{MSE}(\nabla\hat{h}_{s,y},\nabla h_{s,y}),\end{split}$ (2)
where $\hat{p}_{s}$, $\hat{h}_{s,x}$, and $\hat{h}_{s,y}$ represent the predicted foreground probability, horizontal distance, and vertical distance. CE, Dice, and MSE represent cross-entropy loss, dice loss, and mean square error loss.
During the domain adaptation stage, we jointly train the segmentation and detection networks, and the process is shown in Fig. 3 (c) and Stage 2 of Algorithm 1. First, the well-trained segmentation network is utilized to initialize the target domain segmentation model. And the encoder feature of the segmentation network is denoted as $f_{seg}$. Then, we map the feature $f_{seg}$ into a new feature space by a convolution layer and a fully connected layer and the output feature is denoted as $f_{seg}^{c}$ as follows:
@@
where $FC$ and $Conv$ represent fully connected and convolution layers. Finally, the output feature $f_{seg}^{c}$ is used to compute the consistent feature constraint.
In the detection network, considering single central pixel just represents nuclei position while ignoring nuclei texture, it is hard to train an effective detection network using target label $y_{t}^{p}$ [65]. Thus, we employ an extended Gaussian distribution [18] to model the morphology of nuclei. The extended Gaussian encoding process is shown in Fig. 3 (b) and formulated as follows:
@@
where $D_{i}$ represents the distance between pixel $i$ and nearest point annotation, $r_{1}$ represents initialized nuclei radius, the region between $r_{1}$ and $r_{2}$ is background, $\sigma$ is the discrete degree of Gaussian distribution. In $M_{g,i}$, the pixel corresponding to $-1$ does not participate in the training. Employing the generated encoding, a supervised detection model is trained as follows:
$L_{det}=\frac{1}{|\Omega|}\sum_{i\in\Omega}w_{i}(\hat{q}_{t,i}^{(0)}-M_{g,i})^{2},$ (5)
where $\Omega$ represents the pixel set participating training, i.e. $\Omega=\{i|0<D_{i}\leq r_{2}\}$; $w_{i}$ is the training weight of pixel $i$. In order to pay more attention to the foreground, we set $w_{i}$ equal to $10$ when $D_{i}<r_{1}$; otherwise, $w_{i}$ equals $1$. The extended Gaussian encoding not only encodes tissue texture, but can also learn nuclei shape gradually, assisting in segmentation network optimization.
Similar to the above segmentation network, the encoder output $f_{det}$ of the detection network is mapped into new representation $f_{det}^{c}$, which will be used to calculate CFC with $f_{seg}^{c}$ as shown in Fig. 3 (c) and the feature constraint loss $L_{fea}$ is calculated as follows:
@@
where $dim$ represents the dimension of feature encoding. $f_{seg}^{c}$ and $f_{det}^{c}$ are feature embeddings of segmentation and detection networks. Due to the segmentation model being trained on the source domain dataset, its feature representation is closely related to the source domain. In contrast, the detection model is trained on the target domain, so its features are more relevant to the target domain data. Leveraging the target domain representation of $f_{det}^{c}$, the CFC effectively aligns $f_{seg}^{c}$ with $f_{det}^{c}$, enhancing the domain adaptation process.
@@
11:  Initialize: $I=4$; $e_{t}=20$; Initialize segmentation network with $W_{s}$;
@@
15:        Calculate loss: $L_{\text{det}}+\alpha L_{\text{fea}}+\beta L_{\text{dyn}}$ in Eq. 1 and update network;
@@
17:     Prediction outputs: $\hat{p}_{t}$, $\hat{q}_{t}$;
18:     Update pseudo-label: $M_{pse}$leftarrow CPL($\hat{p}_{t}$, $\hat{q}_{t}$, $y_{t}$);